# Dynamic Knowledge Integration for Evidence-Driven Counter-Argument Generation with Large Language Models

### v1

@@ 5 Results
Correlation with Human Judgments. We would like to stress that manual counter-argument evaluation is a highly subjective and tedious process Wachsmuth et al. (2017); Hua et al. (2019); Chen et al. (2024). Thus, ensuring that our automatic evaluation method correlate with human judgments is crucial. Therefore, we computed the correlation between every evaluation metric on the evaluator sample set of 75 examples, including human evaluation, using the method described in Section 4. The Spearman’s rank correlation coefficients ($\rho$) are reported by Figure 2. First, it can be observed that the three LLM-as-a-Judge methods obtain the highest correlation with human judgments (row marked in red). Thus, while JudgeLM and Prometheus obtain a strong correlation, Claude 3.5 Sonnet is the best method with a $\rho$ of 0.82 (very strong correlation).
@@
Reference-based metrics show a poor correlation with both human judgments and LLM-as-a-Judge evaluation methods. This suggests that reference-based metrics might be inadequate for evaluating counter-argument generation quality, as they fail to capture the essential dimensions established for human evaluation and notably disregard the context of the original argument or claim being countered.
Among the LLM-as-a-Judge methods, the lowest correlation was observed between JudgeLM and Prometheus. During a more detailed analysis, we found that these two judges had the highest disagreement in evaluating Opposition and Persuasiveness. Prometheus tends to be stricter on Opposition while JudgeLM does the same on Persuasiveness. Thus, each model assigned lower scores than the other for those particular dimensions.
Overall rank evaluation. Figure 3 reports the overall rankings obtained by summing all the scores from each dimension that each evaluator provided (as described in Section 4). The first observation is that 3 out of 4 LLM-as-a-Judge agree that Command R+ with external knowledge generates the best counter-arguments. Moreover, according to the manual evaluation, both Command R+ and Mistral-7B-Instruct with external knowledge are ranked at the top. Interestingly, results with respect to the ranking of other models vary, but all four evaluators agree that the human-generated counter-arguments (gold reference) are ranked worst.
Table 2 shows the scores obtained by the reference-based metrics. While Command R+ with external knowledge remains the best model, it is easy to see that the rankings are not aligned with respect to human preferences. Furthermore, and although not directly comparable, the obtained scores are in the range of previous work evaluating counter-argument generation with these metrics Chen et al. (2024).
@@
Taking into account the high correlation of LLM-as-a-Judge methods with human preferences, we also evaluated the four models over the full corpus of 150 arguments and counter-arguments pairs. The results from each LLM-as-a-Judge evaluator can be found in the Appendix B (Figures 5, 6, 7, 8).
@@
Point-wise evaluation. Figure 4 illustrates the average dimension-wise rankings from manual evaluation. We can see that Command R+ excels in Opposition compared to others, whereas both Command R+ and Mistral-7B-Instruct with external knowledge are valued the best on Persuasiveness and Relatedness. Moreover, Mistral-7B-Instruct with external knowledge and Command R+ obtain the highest Factuality scores. Overall, we can conclude that, when external evidence is provided to the LLMs, the overall quality of the counter-argument improves. Nevertheless, the performance of generating counter-arguments based on parametric knowledge is quite high, especially for the larger models (Command R+).
@@ 6 Analysis of Results
To answer RQ3, we checked whether the generated counter-arguments indeed used the provided external knowledge, or whether the retrieved evidence is already incorporated in the parametric knowledge of LLMs.
We used the sentence transformer model ‘gte-base-en-v1.5’ Li et al. (2023); Zhang et al. (2024) to calculate sentence-level cosine similarity between the generated outputs and provided external evidence. Our methodology involved segmenting both the external information and generated counter-arguments into individual sentences for pairwise comparison, followed by ranking the similarity scores in descending order. Based on manual verification, we established a threshold of 70% similarity as indicative of successful external knowledge integration. Our analysis revealed that Command R+ with external knowledge effectively utilized external evidence in 82% of its generated counter-arguments, while Mistral-7B-Instruct with external knowledge demonstrated such integration in 51% of the cases.
@@
The topics of high controversy trigger LLMs to start generation by acknowledging the stance of the input argument which exhausts the space for including the retrieved evidence, even when explicitly prompted not to do so. Moreover, such counter-arguments were evaluated better than the output that starts the alternative perspective straight from the factual examples.
@@ 7 Conclusion
In this study, we examined the effect of incorporating externally dynamically retrieved evidence into LLMs in counter-argument generation using web search as an external knowledge source. Our results on a newly created gold dataset show that, while LLMs with external knowledge improved their counter-argument generation, their reliance on it varies, and in some cases, they generate responses with parametric knowledge which obtained better scores.
@@ 8 Limitations
Finally, the LLM-generated counter-arguments may be affected by potential data contamination, where topics and examples of the arguments used in our experiments may overlap with the training data of the LLMs we used. Investigating task contamination is far from trivial but it should be included in any future work.
@@ Acknowledgments
We would like to acknowledge the funding received by several MCIN/AEI/10.13039/501100011033 projects: (i) DeepKnowledge (PID2021-127777OB-C21) and ERDF A way of making Europe; (ii) EDHIA (PID2022-136522OB-C22); (iii) DeepMinor (CNS2023-144375) and European Union NextGenerationEU/PRTR; (iv) DISARGUE (TED2021-130810B-C21). Anar Yeginbergen’s PhD contract is part of the PRE2022-105620 grant, financed by MCIN/AEI/10.13039/501100011033 and by the FSE+.
@@ Appendix A Example of data
Appendix B Per dimension evaluation ranks
@@
In this appendix we provide evaluation ranks per opposition, factuality, relatedness, specificity, and persuasiveness dimensions for the following LLM-as-a-Judge models: JudgeLM in Figure 5, Prometheus in Figure 6 and Claude in Figure 7. The overall ranks over the whole dataset by LLM-as-a-Judge are shown in Figure 8.
@@
Results, especially those computed using Claude 3.5 Sonnet, align with those obtained in Section 5 with a human-annotated sample of the data. Thus, 8 shows that Command R+ with external knowledge is also ranked first, gold reference, and Mistral-7B-Instruct the worst while the ranks for Command R+ and Mistral-7B-Instruct with external knowledge vary according to the judge.
@@
Appendix C Instruction example for human evaluation.
An example of the instructions provided to the human evaluators in the Prolific platform is shown in Figure 9. Firstly, the argument is presented accompanied by a description of the five dimensions (opposition, relatedness, factuality, specificity, persuasiveness) that are evaluated in the paper. Secondly, each of the counter-arguments is presented and the evaluator has to select a value between 1 for unsatisfactory, 2 for moderately satisfactory and 3 for highly satisfactory in each of the dimensions.

### v2

@@ 5 Results
Correlation with Human Judgments. We would like to stress that manual counter-argument evaluation is a highly subjective and tedious process Wachsmuth et al. (2017); Hua et al. (2019); Chen et al. (2024). Thus, ensuring that our automatic evaluation method correlates with human judgments is crucial. Therefore, we computed the correlation between every evaluation metric on the evaluator sample set of 75 examples, including human evaluation, using the method described in Section 4. The Spearman’s rank correlation coefficients ($\rho$) are reported in Figure 2. First, it can be observed that the three LLM-as-a-Judge methods obtain the highest correlation with human judgments (row marked in red). Thus, while JudgeLM and Prometheus obtain a strong correlation, Claude 3.5 Sonnet is the best method with a $\rho$ of 0.82 (very strong correlation).
@@
Reference-based metrics show a poor correlation with both human judgments and LLM-as-a-Judge evaluation methods. This suggests that reference-based metrics may be inadequate for evaluating the quality of counter-argument generation, as they fail to capture the essential dimensions established for human evaluation and disregard the context of the original argument or claim being countered.
Among the LLM-as-a-Judge methods, the lowest correlation was observed between JudgeLM and Prometheus. Upon a more detailed analysis, we found that these two judges had the highest level of disagreement in evaluating Opposition and Persuasiveness. Prometheus tends to be stricter on Opposition, while JudgeLM does the same on Persuasiveness. Thus, each model assigned lower scores than the other for those particular dimensions.
Overall rank evaluation. Figure 3 reports the overall rankings obtained by summing all the scores from each dimension that each evaluator provided (as described in Section 4). The first observation is that 2 out of 3 LLM-as-a-Judge agree that Command R+ with external knowledge generates the best counter-arguments. Moreover, according to the manual evaluation, both Command R+ and Mistral-7B-Instruct with external knowledge are ranked at the top. Interestingly, results with respect to the ranking of other models vary, but all four evaluators agree that the human-generated counter-arguments (gold reference) are ranked worst.
Table 2 shows the scores obtained by the reference-based metrics. While Command R+ with external knowledge remains the best model, it is easy to see that the rankings are not aligned with respect to human preferences. Furthermore, although not directly comparable, the obtained scores are in the range of previous work evaluating counter-argument generation with these metrics Chen et al. (2024).
@@
Taking into account the high correlation of LLM-as-a-Judge methods with human preferences, we also evaluated the four models over the full corpus of 150 arguments and counter-arguments pairs. The results from each LLM-as-a-Judge evaluator can be found in the Appendix E (Figures 5, 6, 7, 8).
@@
Point-wise evaluation. Figure 4 illustrates the average dimension-wise rankings from manual evaluation. We can see that Command R+ excels in Opposition compared to others, whereas both Command R+ and Mistral-7B-Instruct with external knowledge are valued the best on Persuasiveness and Relatedness. Moreover, Mistral-7B-Instruct with external knowledge and Command R+ obtain the highest Factuality scores. Overall, we can conclude that when external evidence is provided to the LLMs, the overall quality of the counter-argument improves. Nevertheless, the performance of generating counter-arguments based on parametric knowledge is quite high, especially for the larger models (Command R+).
@@ 6 Analysis of Results
To answer RQ3, we checked whether the generated counter-arguments indeed used the provided external knowledge or whether the retrieved evidence is already incorporated in the parametric knowledge of LLMs.
We used the sentence transformer model gte-base-en-v1.5’ Li et al. (2023); Zhang et al. (2024) to calculate sentence-level cosine similarity between the generated outputs and provided external evidence. Our methodology involved segmenting both the external information and generated counter-arguments into individual sentences for pairwise comparison, followed by ranking the similarity scores in descending order. Based on manual verification, we established a threshold of 70% similarity as indicative of successful external knowledge integration. Our analysis revealed that Command R+ with external knowledge effectively utilized external evidence in 82% of its generated counter-arguments, while Mistral-7B-Instruct with external knowledge demonstrated such integration in 51% of the cases.
@@
The topics of high controversy trigger LLMs to start generation by acknowledging the stance of the input argument, which exhausts the space for including the retrieved evidence, even when explicitly prompted not to do so. Moreover, such counter-arguments were evaluated better than the output that starts the alternative perspective straight from the factual examples.
@@ 7 Conclusion
In this study, we examined the effect of incorporating externally dynamically retrieved evidence into LLMs in counter-argument generation using web search as an external knowledge source. Our results on a newly created gold dataset show that, while LLMs with external knowledge improved their counter-argument generation, their reliance on it varies, and in some cases, they generate responses with parametric knowledge that obtained better scores.
@@ 8 Limitations
Finally, the LLM-generated counter-arguments may be affected by potential data contamination, where topics and examples of the arguments used in our experiments may overlap with the training data of the LLMs we used. Investigating task contamination is far from trivial, but it should be included in any future work.
@@ Acknowledgments
We are grateful for the provided API credits by the Cohere For AI Research Grant Program [^6] https://cohere.com/research/grants.
We would also like to acknowledge the funding received by the following MCIN/AEI/10.13039/501100011033 projects: (i) DeepKnowledge (PID2021-127777OB-C21) and ERDF A way of making Europe; (ii) DeepMinor (CNS2023-144375) and European Union NextGenerationEU/PRTR; (iii) Disargue (TED2021-130810B-C21) and European Union NextGenerationEU/ PRTR; (iv) EDHIA (PID2022-136522OB-C22); (v) LOTU (TED2021-130398B-C22) and European Union NextGenerationEU/ PRTR. Anar Yeginbergen’s PhD contract is part of the PRE2022-105620 grant, financed by MCIN/AEI/10.13039/501100011033 and by the FSE+.
@@ Appendix A Example of data
Appendix B Examples of arguments excluded due to subjectivity, irrelevance, or lack of verifiability
if a logical outcome of an action or ideology is undesirable, people will try to find a way to avoid that outcome while still pursuing the action or ideology. They argue that people and societies are generally smart and reasonable enough to know when to stop before reaching an undesirable conclusion. if Napoleon had won at Waterloo and completed his conquest of Europe, the continent would have flourished economically, socially, and politically. Santa remains a beloved and benign figure in popular culture. Eating meat is equivalent to murder, and therefore, meat-eaters feel hypocritical calling out others for immoral behavior. It is not fair to judge others for their moral transgressions, such as racism or rape, while they themselves contribute to animal suffering. One should take ownership of their own problems and not place the burden on others to fix them.
Appendix C Snippet of the final version of the dataset
An example of the final version of the data in Table 4.
Argument Counter-Argument
[Increasing privilege for everyone is not possible without those who are currently benefiting losing some of their advantages], and that [people only pay lip service to values like equality and fairness without being willing to make real sacrifices.] [When faced with the choice, most people will prioritize protecting their own privileged position rather than fighting for equality.] [Relinquishing some privilege can be in one’s self-interest, as it can prevent violence and resentment from those who feel underprivileged.] [Affirmative action and recognizing privilege can lead to a more equal society, where no one has advantages based on skin color, sexuality, or gender.] [A society that promotes universal principles and equality can benefit everyone, including those in positions of power.]
[Small countries in Europe should unite as a single nation, like the US], and [eventually all small countries should become large, united territories.] [The US is a union of states with its own laws and cultures], but [it has a high homicide rate compared to the EU, partly due to clashes between different ethnic and socioeconomic groups.] [Closer economic union between European countries may have advantages and disadvantages], but [a political union may not be beneficial as most countries feel they can govern themselves better than a central European government]. [A separate union for smaller countries could be an option, providing financial leverage and allowing them to maintain their social programs and sovereignty.]
[Food past recommended daily values should be heavily taxed], as [overeating is a major problem that leads to health issues and costs money in healthcare.] [I propose a system where individuals would be encouraged to eat the proper amount of food, as calculated by a doctor’s visit], and [those who choose to eat excessively would be taxed to support the system and fund welfare programs.] [The current food distribution system is flawed because it’s based on wealth, favouring the rich], and [a proposed system to tax food based on consumption would not solve the issue of global hunger], which [is caused by factors like war and poor distribution.] [The world produces enough food to feed everyone], but [over 50% of it is wasted], and [the problem of obesity is not solely caused by food consumption.] Instead of taxing food, [it would be more effective to work on improving the distribution system] and [addressing the root causes of global hunger.]
[Capitalism is the best economic option], as [it is a free market regulated by the state] that [allows individuals to become self-made millionaires.] [A just and non-corrupt state is necessary for a successful capitalist environment. Communism, on the other hand, interferes with individual freewill and kills incentives], while [a laissez-faire system is unattainable due to human nature.] [Government intervention interferes with the free market], which [is contrary to pure capitalism.] [Historically, capitalism has been the best option to incentivize the production of physical goods in a world with abundant natural resources.] However, [in a future where physical work is performed by robots and software only needs to be written once], [capitalism may not be the best option.]
Table 4: Final argument and counter-argument pair that was obtained as described in Section 3. Text in Blue highlights the claim, evidence in Orange, and examples are in Red. Criteria to identify claims and premises based on Sviridova et al. (2024).
Appendix D Prompts used for query generation, web search, generation with and without external knowledge
Table 5 illustrates the prompts used for each setting of the generation step.
Setting Prompt
Query Generation Generate a list of 5 queries for web-search that would help to find information to question the veracity of the given claim and persuade to take the opposing position: {claim}?
Provide only questions and nothing else. The answer should be in JSON.
Web-search Answer the question from the following text: {question}?
1. Find factual information from different media outlets
2. Provide the evidence in a bullet point manner.
3. Do not output anything else.
Generation with external knowledge Given the following context: {qacontext}, provide a succinct counter-argument that refutes the following argument using information from the context: {claim}.
Provide only the answer and nothing else.
Make sure the answer is no longer than 3 sentences.
Generation without external knowledge Generate a succinct counter-argument that refutes the following claim: {claim}?
Provide only the answer and nothing else.
Make sure the answer is no longer than 3 sentences.
Table 5: List of prompts used at every stage of the generation. {claim} is a placeholder for the input claim. {question} is a placeholder for the question. {qacontext} is a placeholder for the retrieved context.
Appendix E Per dimension evaluation ranks
@@
In this appendix, we provide evaluation ranks per opposition, factuality, relatedness, specificity, and persuasiveness dimensions for the following LLM-as-a-Judge models: JudgeLM in Figure 5, Prometheus in Figure 6, and Claude in Figure 7. The overall ranks over the whole dataset by LLM-as-a-Judge are shown in Figure 8.
@@
Results, especially those computed using Claude 3.5 Sonnet, align with those obtained in Section 5 with a human-annotated sample of the data. Thus, 8 shows that Command R+ with external knowledge is also ranked first, gold reference, and Mistral-7B-Instruct the worst, while the ranks for Command R+ and Mistral-7B-Instruct with external knowledge vary according to the judge.
@@
Appendix F Instruction example for human evaluation.
An example of the instructions provided to the human evaluators in the Prolific platform is shown in Figure 9. Firstly, the argument is presented accompanied by a description of the five dimensions (opposition, relatedness, factuality, specificity, persuasiveness) that are evaluated in the paper. Secondly, each of the counter-arguments is presented and the evaluator has to select a value between 1 for unsatisfactory, 2 for moderately satisfactory, and 3 for highly satisfactory in each of the dimensions.